### v1

# Draft Model Knows When to Stop: A Self-Verification Length Policy for Speculative Decoding

Abstract
Speculative Decoding (SD) has become an important technique in accelerating the inference speed of large language models. Conventional SD methods employ a fixed draft length, which ignores the token generation difficulty across tasks. Consequently, in this paper, we address such an issue and introduce SVIP - a difficulty-aware dynamic draft length policy for speculative decoding systems. Based on a theoretical lower bound of draft token acceptance rate and its inference-time approximation, SVIP adaptively determines the lengths of draft sequences based on the entropy of each draft token distribution. Experimental results on mainstream SD benchmarks and frameworks demonstrate the superior performance of SVIP, achieving up to 20% walltime speedup on SpecBench over baseline SD methods and 60% speedup on MT-Bench for long-form generation of up to 8K tokens. Moreover, SVIP is totally training-free and compatible with any existing SD methods that generate draft tokens autoregressively. Experimental results also show that SVIP yields consistent walltime improvement on top of GliDe & CaPE and EAGLE-2.
Draft Model Knows When to Stop:
A Self-Verification Length Policy for Speculative Decoding
Ziyin Zhang1,2thanks: Work done during their internship at Tencent AI Lab. Jiahao Xu2 Tian Liang2 Xingyu Chen1,2∗ Zhiwei He1,2∗ Rui Wang1thanks: Corresponding author.Zhaopeng Tu2 1Shanghai Jiao Tong University 2Tencent AI Lab 1{daenerystargaryen,galaxychen,zwhe.cs,wangrui12}@sjtu.edu.cn 2{jettexu,ttianliang,zptu}@tencent.com
1 Introduction
Speculative decoding (Leviathan et al., 2023; Chen et al., 2023) is a novel technique that markedly enhances the generation wall-time of large language models (LLMs). This approach employs a small and efficient amateur model to draft sequences, while concurrently utilizing a larger and more powerful expert model to verify the drafts. By avoiding the autoregressive generation of each token through the target LLM, speculative decoding achieves improved efficiency while preserving the quality of the output.
Figure 1: The “difficulty” of tokens varies in a sequence, resulting in different numbers of accepted draft tokens at different positions.
Many variants of speculative decoding have been proposed. A line of work focuses on developing a stronger and faster draft model (Li et al., 2024b; Elhoushi et al., 2024; Du et al., 2024). Another line of work contributes to maximizing the acceptance probability of draft tokens (Sun et al., 2023; Li et al., 2024a; Lu et al., 2024). In general, they all tend to maximize the alignment between the draft and target model to further maximize the system acceptance rate.
Though successful, most of these works limit their settings to a fixed draft length, where the draft model always generates a fixed number of tokens (e.g. 4 or 5) in each iteration. Such a setting ignores the fact that some tokens - such as stop words or civilities - in the generation may be easy for the draft model to predict, while others - such as knowledge-intensive or reasoning-intensive tokens - can be much harder, as shown in Figure 1.
Figure 2: The correlation between draft model entropy and draft token acceptance probability (top) and lengths of accepted draft seqeunces (bottom).
To address this issue, in this work we introduce SVIP - Self-VerIfication length Policy, a simple, plug-and-play dynamic draft length policy for speculative decoding systems, which enhances the wall time speedup of these systems by adaptively allowing for longer draft sequences for “simple” tokens (top of Figure 1) and terminating the drafting process early upon encountering “hard” tokens (bottom of Figure 1).
Specifically, we first analyze the acceptance rate of speculative decoding systems and derive a lower bound based on the system’s entropy information. Further empirical analysis suggests that such a bound can be approximated by the entropy of the draft model only, which is naturally available in the drafting process of any auto-regressive draft models. Consequently, we develop SVIP  which controls the length of draft sequences dynamically by determining whether to continue drafting or start verification upon the generation of each draft token.
With extensive experiments across multiple model sizes and generation lengths, we demonstrate the superior performance of SVIP. It yields more than 20% of improvements over vanilla speculative decoding for Qwen2.5 14B and LLaMA-3 70B on SpecBench (Xia et al., 2024), and more than 60% of improvements for Pythia 6.9B on MT-Bench (Zheng et al., 2023) when generating long-form responses of up to 8K tokens.
Moreover, since our method is lightweight and training-free, it is extremely flexible and can be adapted to any speculative decoding system with an auto-regressive draft model. As examples, we apply SVIP on top of two state-of-the-art speculative decoding methods: GliDe with a CaPE (Du et al., 2024) and EAGLE-2 (Li et al., 2024a), and confirm that it brings consistent improvements. Our code is available at https://github.com/Geralt-Targaryen/SVIP.
In summary, our contributions are threefold:
  1. 1.
    We derive a low bound of speculative decoding systems, where the acceptance rate of the draft model could be modeled by its entropy only.
  2. 2.
    Based on this lower bound, we further develop an entropy-based dynamic draft length policy for speculative decoding systems, which is extremely flexible and can be adapted to any auto-regressive draft model.
  3. 3.
    Experimental results demonstrate the superior performance of SVIP over baseline methods, with up to 20% average speedup on SpecBench, 60% in long-form generation, and consistent improvement over state-of-the-art speculative decoding frameworks such as GliDe & CaPE and EAGLE-2.
2 Method
Figure 3: Distribution histograms of the entropy ratio $H_{q,p}/H_{q}$. For most tokens, this ratio falls into a narrow range, indicating that the cross entropy $H_{q,p}$ can be approximated by a constant multiplication of the draft entropy $H_{q}$.
The overall objective of SVIP is to dynamically adapt draft length on-the-fly, stopping early if the current draft token’s acceptance probability is low and otherwise continuing drafting. To introduce our method, we first provide the background on speculative decoding in Section 2.1. Then, based on the key observation that the acceptance probability of a draft token depends on the target model confidence - which is unavailable in the drafting phase, we derive a theoretical lower bound for the acceptance rate based on the draft model’s entropy and the cross-entropy between the target and draft models in Section 2.2. Finally, based on empirical analysis of the target and draft distributions of more than 100K tokens across three model families, we approximate this lower bound using only the entropy information of the draft model in Section 2.3, making it viable in actual inference.
2.1 Preliminaries on Speculative Decoding
Suppose we have two LLMs $p$ and $q$, where $p$ is the larger (target) model, and $q$ is the smaller (draft) model. Given an input sequence $x_{\leqslant t}$ of length $t$, and a draft length $\gamma$, the draft model first samples $\gamma$ tokens $x_{t+1},\cdots,x_{t+\gamma}$ autoregressive, which are verified by the target models in parallel to acquire the confidences $p(x_{t+1}),\cdots,p(x_{t+\gamma})$ [^1] The short hand $p(x_{n})$ is used to denote the conditional probability $p(x_{n}|x_{<n})$ when there is no ambiguity. Throughout the work we use subscripts to indicate token indices in a sequence (e.g. $x_{n}$ for $n$-th token), and superscripts to indicate element indices in a vector (e.g. $x_{n}^{i}$ for $i$-th element in $x_{n}$)..
[…]
The complete algorithms for speculative decoding are given in Appendix A.
Figure 4: Comparison between the actual acceptance probability from Equation (2), the acceptance probability lower bound from Equation (5), and the estimated lower bound after approximating the cross entropy $H_{q,p}$ with a constant multiplication of $H_{q}$. Each position on the x-axis corresponds to a token, which has been sorted according to the actual acceptance probability.
2.2 Theoretical Lower Bound of Acceptance Rate
From Section 2.1, it’s easy to derive that given an input sequence $x_{<t}$ and a draft token $x_{t}$, its acceptance probability is $\min\left(1,\;\frac{p(x_{t})}{q(x_{t})}\right)$. Let $\beta$ denote the expected acceptance probability over the distribution of $x_{t}$, and it follows that
$\beta$ $=\sum_{x}q(x)\cdot\min\left(1,\frac{p(x)}{q(x)}\right)$ (1)
$=\sum_{x}\min\left(p(x),q(x)\right).$ (2)
Chen et al. (2023) has proven that $\beta$ is related to the total variational distance (TVD) between $p$ and $q$:
$\beta=1-\mathrm{TVD}(p,q).$ (3)
According to Pinsker’s inequality - which relates TVD to Kullback-Leibler divergence - we then have
$\beta$ $\geqslant 1-\sqrt{\frac{1}{2}\mathrm{KL}(q||p)}$ (4)
$=1-\sqrt{\frac{1}{2}\sum_{x}q(x)\log\frac{q(x)}{p(x)}}$
$=1-\sqrt{\frac{1}{2}H_{q,p}-\frac{1}{2}H_{q}},$ (5)
where $H_{q,p}$ is the cross entropy between $q$ and $p$, and $H_{q}$ is the entropy of $q$.
Equation (5) provides a theoretical lower bound for the acceptance rate of a draft token using the entropy information of the speculative decoding system [^2] An alternative way to construct the lower bound is discussed in Appendix B.. To provide an intuitive motivation for using entropy to construct the lower bound, we plot the relation between draft model entropy and draft token acceptance probability as well as accepted draft seqeunce lengths in Figure 2, which are collected from more than 1M tokens generated by three different target models with temperature set to 1 and maximum draft length set to 40. The first subfigure indicates a strong negative correlation between draft model entropy and draft token acceptance probability, while the second subfigure shows that when the entropy is low, dozens of consecutive draft tokens could be accepted, which highlights the drawback of setting draft length to a constant, small value, as is the common practice in speculative decoding literature. Another version of Figure 2 in the greedy setting is given in Appendix D.
2.3 Empirical Estimation of Acceptance Rate
So far in this section, we have been assuming access to the target distribution $p(x_{t})$ of the next token, which is unavailable in the drafting phase at inference time. Thus, to apply Equation (5) for actual acceptance rate estimation, we must approximate the cross entropy $H_{q,p}$ with information from only the draft model.
Algorithm 1 SVIP
1:target model $p$, draft model $q$, input sequence $x_{\leqslant t}$, maximum length $T$, threshold $h$
2:Initialize $n\leftarrow t$
3:while $n<T$ do
4:     $j=0$
5:     while True do
6:         Sample $x_{n+j}\sim q(x|x_{<{n+j}})$
7:         $j\leftarrow j+1$
8:         if $\sqrt{H(q_{x|x_{<n+j}})}>h$ then
9:              Exit while loop
10:         end if
11:     end while
12:     $\gamma\leftarrow j$
13:     Compute $p(x|x_{<{n+j}}),\;j=1,\cdots,\gamma+1$ in parallel
14:     $\tilde{n}\leftarrow n$
15:     for $j=1$ to $\gamma$ do
16:         if Verify$\left(p_{x|x_{<n+j}},\,q_{x|x_{<n+j}},x_{n+j}\right)$ then
17:              $\tilde{n}\leftarrow\tilde{n}+1$
18:         else
19:              $x_{n+j}\leftarrow\text{Correct}\left(p_{x|x_{<n+j}},\,q_{x|x_{<n+j}}\right)$
20:              Exit for loop
21:         end if
22:     end for
23:     if $\tilde{n}==n+\gamma$ then
24:         Sample $x_{n+\gamma+1}$ from $p(x|x_{\leqslant n+\gamma})$
25:     end if
26:     $n\leftarrow\tilde{n}+1$
27:end while
28:$x_{\leqslant n}$
To tackle this issue, we first plot the relationship between $H_{q,p}$ and $H_{q}$ in Figure 3. For all three model families, the entropy ratio $H_{q,p}/H_{q}$ are concentrated in a narrow range between 1.0 and 1.3. Thus, we choose to approximate $H_{q,p}$ with a constant multiplication of $H_{q}$. Plugging it into Equation (5), we now have
$\beta$ $\geqslant 1-\sqrt{cH_{q}},$ (6)
where $c$ is a constant controlling the approximation ratio between $H_{q,p}$ and $H_{q}$. In Figure 4, we visualize the values derived from Equation (2), (5), and (6).
With Equation (6) providing a way to estimate acceptance probability using only the draft model’s entropy, we can now adapt the draft length on-the-fly. After generating each draft token, we compute the estimated acceptance probability lower bound, and stop the draft process if it’s lower than a certain threshold $h$. We note that since both $c$ and $h$ are constant hyperparameters, we can remove $\sqrt{c}$ from Equation 6 and absorb it into the threshold $h$.
We formalize SVIP in Algorithm 1. The details of the methods Verify and Correct are given in Appendix A, for which different versions are available for sampling (Algorithm 2, 4) and greedy decoding (Algorithm 3, 5).
3 Experiments
3.1 Experiments on SpecBench
3.1.1 Settings
Methods MT-Bench Trans. Sum. QA Math RAG Avg.
Pythia (6.9B, ${}_{\text{(6.9B, 160M)}}$ Const. 1.45 1.47 1.24 1.43 1.52 1.42 1.42
Heuristics 1.51 1.58 1.34 1.58 1.64 1.51 1.53
SVIP 1.63 1.62 1.45 1.67 1.72 1.66 1.63(+14.8%)
Qwen2.5 (14B, ${}_{\text{(14B, 0.5B)}}$ Const. 1.08 0.87 1.11 0.92 1.43 0.99 1.07
Heuristics 1.10 0.91 1.10 0.92 1.34 1.03 1.07
SVIP 1.33 1.12 1.37 1.14 1.57 1.23 1.29(+20.6%)
LLaMA-3 (70B, ${}_{\text{(70B, 8B)}}$ Const. 2.04 2.48 2.56 2.34 2.32 2.28 2.34
Heuristics 2.30 3.13 3.33 2.61 2.52 2.63 2.76
SVIP 2.31 3.04 3.48 2.63 2.89 2.59 2.83(+20.9%)
Table 1: Speedup over target-model-only autoregressive decoding on SpecBench.
We validate the effectiveness of SVIP on Spec-Bench (Xia et al., 2024) using three distinct target models: Pythia 6.9B (Biderman et al., 2023), Qwen2.5 14B (Yang et al., 2024), and LLaMA-3 70B (Dubey et al., 2024), with Pythia-160M, Qwen2.5 0.5B, and LLaMA-3 8B as the draft models respectively.
As baselines, we consider two simple policies for draft length: 1) a constant draft length of 5, which is commonly used in the literature, and 2) the heuristics implemented in Hugging Face Transformers library (Wolf et al., 2019), where the draft length for the next draft iteration is increased by 2 if all draft tokens in the current iteration are accepted, and otherwise decreased by 1.
We set the sampling temperature to 0 on SpecBench (the alternatives are discussed in Appendix C). For each model, the entropy threshold $t$ in SVIP is chosen from $\{0.2,0.3,0.4,0.5\}$ based on performance on 8 samples held out from MT-Bench (Zheng et al., 2023), which are 0.4 for Pythia, and 0.3 for Qwen2.5 and LLaMA-3. All experiments with Pythia and Qwen are conducted on a single 40GB A100, while experiments with LLaMA are conducted on 5 40GB A100s. To mitigate the impact of system performance varitations, we repeat all experiments with Pythia and Qwen for three times (using different random seeds when they are used) and report the average speedup over target-model-only autoregressive decoding. Also, since the memory consumption of verifying $n$ draft tokens is quadratic in $n$, we limit the maximum draft length to 40 in both heuristics and SVIP scenarios, beyond which we start to encounter out-of-memory issues.
3.1.2 Results
The results on SpecBench are shown in Table 1. Compared with the constant approach, SVIP yields an average speedup of 15% for Pythia and 20% for Qwen and LLaMA. Compared with the heuristic approach, SVIP also gives a consistent improvement on all domains for Pythia and Qwen, and outperforms the latter on 4 out of 6 domains for LLaMA.
Figure 5: The average generated draft length, accepted draft length, and acceptance rate of Qwen2.5 (top) and LLaMA-3 (bottom) on SpecBench. Compared with the two baselines, SVIP leads to a shorter draft length and a much higher acceptance rate.
In Figure 5, we plot the average draft length and accepted draft length of Qwen and LLaMA (the results for Pythia are similar, and are given in Appendix C). From the figure, we observe that by terminating the draft process when the draft model entropy is high, SVIP leads to shorter draft lengths and a much higher acceptance rate (close to or more than 80% on all domains for Qwen, and more than 90% for LLaMA). Notably, the acceptance rate of LLaMA-3 even reaches 99% on the summarization domain, which contributes to the highest speedup (3.48) in Table 1.
3.2 Long-form Generation
Most existing works on speculative decoding (Chen et al., 2023; Du et al., 2024) limit their experiments to generating short sequences of 128 tokens. To verify the wide applicability of SVIP, we also conduct experiments on long-form generation with up to 8K context. For this purpose, we use MT-Bench (Zheng et al., 2023) as the dataset and set the sampling temperature to 1, as we found that when using greedy decoding in long-form generation, both the draft and the target models are prone to repeat themselves, resulting in very low information entropy (see Appendix D for details). Other settings follow Section 3.1.
Methods Generation Length
128 256 512 1K 2K 4K 6K 8K
Pythia (6.9B, ${}_{\text{(6.9B, 160M)}}$ Const. 0.68 0.69 0.69 0.70 0.72 0.90 0.89 0.87
Heuristics 0.88 0.88 0.88 0.88 0.90 1.21 1.25 1.23
SVIP 1.07 1.08 1.08 1.07 1.08 1.43 1.44 1.41
Qwen2.5 (14B, ${}_{\text{(14B, 0.5B)}}$ Const. 0.98 0.97 0.95 0.96 0.98 1.00 1.02 1.04
Heuristics 1.01 0.99 0.98 1.00 1.02 1.03 1.04 1.06
SVIP 1.29 1.29 1.30 1.31 1.32 1.33 1.33 1.35
LLaMA-3 (70B, ${}_{\text{(70B, 8B)}}$ Const. 1.74 1.74 1.77 1.78 1.82 1.90 1.93 1.94
Heuristics 1.53 1.56 1.61 1.63 1.68 1.77 1.83 1.84
SVIP 1.69 1.72 1.75 1.78 1.86 1.96 2.01 2.02
Table 2: Speedup on MT-Bench with different generation length.
The results are given in Table 2. Interestingly, we find that in the sampling setting, the two baseline methods (constant and heuristics) perform even worse than target-model-only auto-regressive decoding for Pythia and Qwen, while SVIP consistently yields a positive speedup. For LLaMA, while constant draft length performs better with contexts shorter than 1K tokens, SVIP exceeds it for generating longer sequences. Another observation is that the speedup ratio of all three methods generally increases with the context length, which could be possibly attributed to the longer contexts giving the draft model more information, aligning it at test time to the target model. However, we note that the absolute values of token throughput (which are not discussed in this paper, since they heavily depend on the underlying machines) stay at the same level from 1K to 8K context for Qwen2.5 and LLaMA-3 when using speculative decoding, and decrease slowly when not using speculative decoding. For Pythia, which does not use any optimized attention such as MQA (Shazeer, 2019) or GQA (Shazeer, 2019), the throughput decreases notably with context length in both cases - with or without specualtive decoding.
3.3 Applying SVIP to Other Draft Methods
In Section 3.1 and 3.2, we evaluated SVIP on vanilla speculative decoding, where a standard pretrained Transformer decoder model from the target model’s family is used as the draft model. However, in the past years many works on speculative decoding have proposed other stronger or more efficient draft models (Cai et al., 2024; Du et al., 2024; Li et al., 2024b). Since most of these works assume a constant draft length, SVIP is orthogonal to them and can be applied on top of them without any additional training.
Methods MT-Bench Code Finance GSM Spider Avg.
7B GliDe 1.95 2.04 1.91 1.98 1.69 1.95
   +SVIP 2.00 2.12 2.03 2.01 1.63 2.02
GliDe + CaPE 2.36 2.57 2.29 2.51 1.97 2.40
   +SVIP 2.56 2.65 2.49 2.54 2.08 2.52
13B GliDe 2.22 2.41 2.15 2.31 1.85 2.24
   +SVIP 2.31 2.43 2.17 2.35 1.85 2.28
GliDe + CaPE 2.73 2.86 2.66 2.80 2.24 2.73
   +SVIP 2.72 2.93 2.66 2.85 2.27 2.76
33B GliDe 2.12 2.25 2.09 2.29 1.99 2.18
   +SVIP 2.29 2.40 2.20 2.42 2.03 2.30
GliDe + CaPE 2.08 1.98 2.10 2.13 1.76 2.03
   +SVIP 2.13 2.02 2.15 2.16 1.82 2.08
Table 3: Speedup comparison with GliDe & CaPE, using Vicuna as the base model.
Specifically, we consider GliDe with a CaPE (Du et al., 2024), where the draft model - named GliDe - is a small transformer decoder with cross-attention to the target model’s hidden representations, while CaPE is a complementary tree expansion method to increase the acceptance rate of the draft token at each position. Following the settings of Du et al. (2024), we use Vicuna 7B, 13B, and 33B (Chiang et al., 2023) as the base models (for which the draft models are publicly available) and set the sampling temperature to 0. However, we distribute the 33B model across two GPUs due to memory constraint. Similar to the previous experiments, we set the threshold $t$ to 0.5 based on pilot experiments on 8 samples held out from MT-Bench using GliDe only (without CaPE).
We also apply SVIP to EAGLE-2 (Li et al., 2024a), the state-of-the-art speculative decoding system which utilizes the target model’s language modeling head on top of the draft model’s features to predict the next draft token, and dynamically constructs a draft tree at each draft position. Following Li et al. (2024a), we use LLaMA-2 7B, 13B (Touvron et al., 2023) and Vicuna 7B, 13B (Chiang et al., 2023) as the base models, and set the sampling temperature to 1. As Brown et al. (2024) suggest, adding a conditional clause (that decides whether or not to stop drafting) after the generation of each draft token in EAGLE-2 may interrupt the otherwise input-independent control flow of EALGE-2, reducing the effects of low-level interpreter and system optimizations. Thus, in EAGLE-2 we calculate the draft entropy and decide whether or not to stop drafting after every two draft tokens instead of after every draft token, as used in previous experiments.
Methods MT-Bench H-Eval GSM8K Alpaca CNN/DM QA Avg.
LLaMA-2 7B EAGLE-2 3.10 3.61 3.15 3.10 2.76 2.84 3.10
   + SVIP 3.16 3.66 3.18 3.13 2.82 3.02 3.16
LLaMA-2 13B EAGLE-2 3.38 4.12 3.41 3.25 3.01 2.98 3.41
   + SVIP 3.41 4.09 3.45 3.34 3.05 3.17 3.46
Vicuna 7B EAGLE-2 2.66 2.84 2.77 2.48 2.31 2.13 2.58
   + SVIP 2.84 2.97 2.75 2.66 2.42 2.30 2.73
Vicuna 13B EAGLE-2 2.85 3.31 2.93 2.74 2.48 2.30 2.83
   + SVIP 2.94 3.49 3.19 2.89 2.60 2.53 2.98
Table 4: Speedup comparison with EAGLE-2, using LLaMA-2-Chat and Vicuna as the base models.
The results for these two methods are presented in Table 3 and 4, respectively. For GliDe, SVIP yields a consistent speedup both with and without CaPE, with a 5% improvement for the 7B and 33B models. For EAGLE-2, the speedup is also particularly notable for the Vicuna models, with 5% improvement for both the 7B and 13B models.
4 Related Work
Since Leviathan et al. (2023) and Chen et al. (2023) introduced speculative decoding into large language models, numerous works have followed their tracks in pursuit of more efficient LLM inference. We broadly categorize these works into three types: better draft models, draft tree expansion, and draft length control, which are orthogonal to each other. A more comprehensive review of speculative decoding is provided by Xia et al. (2024).
Better draft models.
As Xia et al. (2024) suggest, draft models in speculative decoding can be either based on self-drafting or based on an independent draft model. For the first type, one may use a quantized (Zhao et al., 2024), early-exiting (Elhoushi et al., 2024), or forward-padded (Monea et al., 2023) version of the target model to produce draft tokens, while the second type is represented by the vanilla speculative decoding (Leviathan et al., 2023). Some works also take the best of both worlds and introduce extra layers on top of the target model’s hidden representations to construct draft models, represented by EAGLE (Li et al., 2024b), GliDe (Du et al., 2024), and Medusa (Cai et al., 2024).
Draft tree expansion.
Given a draft model, one may verify multiple draft tokens for the same position in parallel to increase the probability of finding an accepted draft token, and we use “draft tree expansion” as an umbrella term for such techniques. Li et al. (2024a) introduce EAGLE-2, which reranks draft tokens in EAGLE’s draft tree to select tokens with the highest confidence for verification. Similarly, CaPE (Du et al., 2024) improves GliDe by expanding the token set chosen for verification at each position based on top-1 confidence. Other works have also addressed the problem of multi-draft verification from a theoretic perspective (Sun et al., 2023; Yin et al., 2024).
Draft length control.
Works in this category are few, but most relevant to ours. Liu et al. (2024) introduce PEARL, which lets the target model perform verification in parallel to draft generation, stopping the draft process when a mismatch is found. Huang et al. (2024) propose SpecDec++, which trains an acceptance prediction head on top of the draft model to predict the acceptance probability of the current draft token, stopping the draft round when the predicted acceptance probability falls below a constant threshold. Brown et al. (2024) propose Dynamic Depth Decoding on top of EAGLE-2, which uses the sum of all tokens’ confidences in one level of its draft tree as an indicator to predict whether or not to continue draft generation.
5 Conclusion
We propose SVIP, a flexible, training-free, and plug-and-play dynamic draft length policy for speculative decoding systems. Based on a theoretical lower bound of acceptance probability and its empirical approximation, SVIP determines whether to continue draft generation or to quit drafting based on the draft model’s entropy after the generation of each draft token. With extensive experiments spanning various base models, draft methods, test domains, and generation length, we validated the effectiveness of SVIP, sparking new insights on speculative decoding and more efficient large language models.
Appendix A The Complete Speculative Decoding Algorithms
In Algorithm 2 to 6, we present the complete algorithms of the vanilla speculative decoding in both the greedy decoding and the sampling scenarios. For the sampling scenario, the Verify and Correct methods in Algorithm 6 resolve to Algorithm 2 and 4. For greedy decoding, they resolve to Algorithm 3 and 5.
Algorithm 2 Verify (Sampling)
1:target distribution $p(x)$, draft distribution $q(x)$, draft token $x_{t}$
2:$accept\leftarrow$ False
3:$r\sim U[0,1]$
4:if $r<\frac{p(x_{t})}{q(x_{t})}$ then
5:     $accept\leftarrow$ True
6:end if
7:$accept$
Algorithm 3 Verify (Greedy)
1:target distribution $p(x)$, draft distribution $q(x)$, draft token $x_{t}$
2:$accept\leftarrow$ False
3:if $\arg\max p(x)==x_{t}$ then
4:     $accept\leftarrow$ True
5:end if
6:$accept$
Algorithm 4 Correct (Sampling)
1:target distribution $p(x)$, draft distribution $q(x)$
[…]
3:$\hat{x}$
Algorithm 5 Correct (Greedy)
1:target distribution $p(x)$, draft distribution $q(x)$
2:$\arg\max p(x)$
Algorithm 6 Speculative Decoding
1:target model $p$, draft model $q$, input sequence $x_{\leqslant t}$, maximum length $T$, draft length $\gamma$
2:Initialize $n\leftarrow t$
3:while $n<T$ do
4:     for $j=1$ to $\gamma$ do
5:         Sample $x_{n+j}\sim q(x|x_{<{n+j}})$
6:     end for
7:     Compute $p(x|x_{<{n+j}}),\;j=1,\cdots,\gamma+1$ in parallel
8:     $\tilde{n}\leftarrow n$
9:     for $j=1$ to $\gamma$ do
10:         if Verify$\left(p(x|x_{<n+j}),\,q(x|x_{<n+j}),x_{n+j}\right)$ then
11:              $\tilde{n}\leftarrow\tilde{n}+1$
12:         else
13:              $x_{n+j}\leftarrow\text{Correct}\left(p(x|x_{<n+j}),\,q(x|x_{<n+j})\right)$
14:              Exit for loop
15:         end if
16:     end for
17:     if $\tilde{n}==n+\gamma$ then
18:         $x_{n+\gamma+1}\sim p(x|x_{\leqslant n+\gamma})$
19:     end if
20:     $n\leftarrow\tilde{n}+1$
21:end while
22:$x_{\leqslant n}$
Appendix B Alternatives for Acceptance Rate Lower Bound Computation
In Section 2, we used Pinsker’s inequality to compute a lower bound for the expected acceptance probability:
$\beta$ $=\sum_{x}\min\left(p(x),q(x)\right)$ (7)
$\geqslant 1-\sqrt{\frac{1}{2}KL(q||p)}.$ (8)
Another way to compute the lower bound of acceptance probability can be derived from Bretagnolle-Huber inequality (Bretagnolle and Huber, 1978):
$\beta$ $\geqslant 1-\sqrt{1-e^{-KL(q||p)}}.$ (9)
Compared with the Pinsker’s bound, it’s trivial to see that this bound is guaranteed to be always larger than 0. However, in practice we find that the Pinsker’s bound is 11% tighter for Qwen2.5, 20% tighter for Pythia, and 43% tighter for LLaMA-3.
Appendix C Additional Results on SpecBench
Methods MT-Bench Trans. Sum. QA Math RAG Avg.
[…]
Heuristics 0.82 0.83 0.85 0.83 0.83 0.83 0.83
SVIP 1.05 1.02 1.03 1.01 1.03 1.00 1.02(+56.9%)
[…]
Heuristics 1.02 0.94 0.93 0.88 1.22 0.91 0.99
SVIP 1.24 1.08 1.19 1.11 1.47 1.10 1.20(+25.0%)
[…]
Heuristics 1.56 1.55 1.76 1.49 1.61 1.55 1.58
SVIP 1.53 1.53 1.69 1.51 1.71 1.56 1.58(-1.3%)
Table 5: Speedup on SpecBench using temperature sampling.
In Figure 6, we plot the draft length and acceptance rate of Pythia (which complements Figure 5 in Section 3.1), and in Figure 7 we also give the results of the three models when using sampling instead of greedy decoding [^3] We note that in the “constant” scenario, the draft length is always set to 5. However, if an EOS token is sampled from the draft model, the draft process will terminate. So the overall draft length might be slightly lower than 5, as can be seen from Figure 7.. Across the different models and sampling methods, the same observation as discussed in Section 3 holds: SVIP results in short draft lengths and a much higher acceptance rate.
Figure 6: The average generated draft length, accepted draft length, and acceptance rate of Pythia on SpecBench.
Figure 7: The average generated draft length, accepted draft length, and acceptance rate of Qwen2.5 (top), Pythia (middle), and LLaMA-3 (bottom) on SpecBench, using temperature sampling instead of greedy decoding.
Appendix D Additional Results on Long-form Generation
Methods Generation Length
128 256 512 1K 2K 4K 6K 8K
Pythia (6.9B, ${}_{\text{(6.9B, 160M)}}$ Const. 1.10 1.30 1.50 1.66 1.82 1.20 1.11 1.04
Heuristics 1.25 1.45 1.65 1.81 2.02 1.50 1.46 1.41
SVIP 1.41 1.62 1.83 2.01 2.21 1.71 1.60 1.52
Qwen2.5 (14B, ${}_{\text{(14B, 0.5B)}}$ Const. 1.05 1.08 1.15 1.29 1.44 1.54 1.60 1.67
Heuristics 1.04 1.06 1.13 1.32 1.54 1.72 1.85 1.97
SVIP 1.30 1.34 1.42 1.57 1.74 1.87 1.98 2.10
LLaMA-3 (70B, ${}_{\text{(70B, 8B)}}$ Const. 2.06 2.18 2.31 2.45 2.58 2.72 2.77 2.78
Heuristics 2.26 2.46 2.73 3.07 3.48 3.90 4.15 4.26
SVIP 2.31 2.56 2.86 3.21 3.59 4.00 4.23 4.33
Table 6: Speedup on MT-Bench with different generation length using greedy decoding.
Figure 8: The correlation between draft model entropy and lengths of accepted draft seqeunces in the greedy setting.
In Table 6, we present the results of long-form generation with greedy decoding. Compared with Table 2, the speedup ratio of greedy decoding is much higher (even more than 4 times for LLaMA-3 after 4K context). This is due to the fact that these models tend to repeat themselves in greedy long-form generation, making it very easy for the draft models to predict the next tokens. In Figure 8, we also plot the relation between draft model entropy and accepted draft sequence lengths in this setting, which shows a much stronger correlation compared with the sampling setting in Figure 2.